Please explain how to write the ff. rational number as the quotient of two integers in simplest form.
_
a.0.372
__
b.0.63

Answer by stanbon(48565)

a.0.372
__= 372/1000= 186/500 = 93/250
b.0.63 = 63/100
Cheers,
Stan H.